OpenStudy (foxythepirate0984):

Valerie plays a game using the spinner shown. During the game she will spin a total of 36 times. Which is the best prediction for the number of times she will get either "lose a turn" or "back 3" on the spinner? A. 2 B. 6 C. 12 D. 18

OpenStudy (foxythepirate0984):

@ChantySquirrel1129** @prettypink_mega @shamim

jimthompson5910 (jim_thompson5910):

how many spaces are labeled either "lose a turn" or "back 3" ?

OpenStudy (foxythepirate0984):

?

jimthompson5910 (jim_thompson5910):

look at the spinner

OpenStudy (foxythepirate0984):

k

jimthompson5910 (jim_thompson5910):

how many times does "lose a turn" come up?

OpenStudy (foxythepirate0984):

1

jimthompson5910 (jim_thompson5910):

how about "back 3"

OpenStudy (foxythepirate0984):

1

jimthompson5910 (jim_thompson5910):

so there are 2 spaces out of 6 2/6 = 1/3

jimthompson5910 (jim_thompson5910):

1/3 is the probability of landing on either "lose a turn" or "back 3"

jimthompson5910 (jim_thompson5910):

multiply that probability with the number of spins (36)

OpenStudy (foxythepirate0984):

1/3 x 36 = 3?

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (foxythepirate0984):

16?

jimthompson5910 (jim_thompson5910):

it might help to think of 1/3 as 0.3333 (approximate)

jimthompson5910 (jim_thompson5910):

1/3 times 36 = 0.3333 times 36 = ???

OpenStudy (foxythepirate0984):

11.9988

jimthompson5910 (jim_thompson5910):

round that to the nearest whole number

OpenStudy (foxythepirate0984):

umm 19?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (foxythepirate0984):

18

jimthompson5910 (jim_thompson5910):

I would appreciate it if you didn't randomly guess

jimthompson5910 (jim_thompson5910):

11.9988 has the whole part of "11" the number 11.9988 is between 11 and 12

OpenStudy (foxythepirate0984):

12?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (foxythepirate0984):

Thx for the help

jimthompson5910 (jim_thompson5910):

yw